### v3

# A proof of the reverse isoperimetric inequality using a geometric–analytic approach

Abstract
We present a proof of the reverse isoperimetric inequality—a central conjecture in extended black hole thermodynamics—for black holes in Einstein gravity with $D\geq 4$, employing a two-pronged geometric–analytic method. Our analysis shows that the reversal of the usual isoperimetric inequality originates from the structure of curved backgrounds governed by Einstein’s equations, thereby underscoring the fundamental role of gravity in the reverse isoperimetric property of AdS black hole horizons.
I Introduction
Extended black hole thermodynamics [17, 8, 5] provides a richer structure of black hole thermodynamics by identifying the cosmological constant $\Lambda$ with pressure as
$P=-\frac{\Lambda}{8\pi}$ (1)
This is possible for the AdS case for which we have $\Lambda<0$, which implies that $P>0$ as required on physical grounds. The conjugate thermodynamic potential is given as
$\Theta=-\frac{(D-2)}{16\pi(D-1)}r_{h}A-\frac{1}{2(D-1)}\sum_{i}a_{i}J_{i},$ (2)
where $a_{i}$ are the rotation parameters, $J_{i}$ are the angular momenta of the black hole, and $A$ is the area of outer horizon. This leads to the modified first law with a varying cosmological constant as
$dM=TdS+\Omega_{i}dJ_{i}+\Theta d\Lambda.$ (3)
The modified first law points to the mass of the AdS black hole as being the enthalpy of spacetime [17]. This has given rise to interesting phenomena concerning black holes, such as Van der Waals fluids [2, 18], and heat engines [15].
A varying cosmological constant has been shown to arise from higher-dimensional bulk effects, in which case the varying brane tension $\tau$ induces extended thermodynamics on the brane [11]. The modified first law has been shown to arise robustly via the Extended Iyer-Wald formalism in [24]. In the context of extended phase space, an interesting result is conjectured originally within Einstein’s gravity called the “reverse” isoperimetric inequality (RII) [5], which is formally defined as
$\Bigg(\frac{(D-1)V}{\mathcal{A}_{D-2}}\Bigg)^{\frac{1}{D-1}}\geq\Bigg(\frac{A}{\mathcal{A}_{D-2}}\Bigg)^{\frac{1}{D-2}},$ (4)
where $V$ is the thermodynamic volume, $\mathcal{A}_{D-2}$ is the volume of unit sphere and $A$ is the area of the outer horizon. Remarkably, for AdS–Schwarzschild black holes, the definition of thermodynamic volume equals the geometric volume defined as the integral [5]
$V_{\text{geo}}=V^{\prime}=\int_{r_{0}}^{r_{h}}dr\int d\Omega\sqrt{-g},$ (5)
which is over the interior of the black hole, where we vary the radial coordinate from the singularity at $r=r_{0}$ to the outer horizon at $r=r_{h}$. In fact, the thermodynamic and geometric volumes are the same for all static, spherically symmetric black holes. For Kerr–AdS, the geometric volume obeys the standard isoperimetric inequality while the thermodynamic volume obeys the reverse isoperimetric inequality. Therefore, Cvetic et al. stated the conjecture as—At a fixed thermodynamic volume $V$ , the black hole with the largest entropy is Schwarzschild-AdS. Furthermore, the geometric volume for the AdS–Schwarzschild black hole evaluates to
$V^{\prime}=\frac{4}{3}\pi r_{h}^{3}$ (6)
as if it were just the Euclidean volume $V_{E}$ inside a (D2)(D-2)-sphere in Euclidean space $\mathbb{E}^{D-1}$. In this case, we specifically have
$V_{\text{thermo}}=V_{\text{geo}}=V_{E}.$ (7)
This leads us to interpret the conjecture as the statement that a round sphere $S^{D-2}$ in $D$ dimensional AdS space, which bounds a fixed geometric volume equal to the volume of a (D1)(D-1) Euclidean ball, maximizes entropy. Physically, this means that black holes like to be round [1]. Therefore, although the relevant notion of volume is “thermodynamic volume” in the extended black hole thermodynamics, in the static case of Schwarzschild-AdS black hole, this coincides with the naive geometric volume as defined in (5), and we shall express calculations in that language with the understanding that they represent thermodynamic volume in the static case.
The inequality is reversed in the sense that in Euclidean space, a round sphere minimizes area, called simply the isoperimetric inequality [21, 22]. The reverse isoperimetric inequality is known to be obeyed for every case except for the charged Bañados-Teitelboim-Zanelli (BTZ) black holes [19, 10]. Black holes that violate RII are called superentropic [14, 13]. However, these are thermodynamically unstable [16] since they have negative heat capacity at constant volume. RII is a classical result, and some quantum inequalities have also been proposed recently regarding this [9]. Despite the success of RII, it lacks a general proof and remains a conjecture.
In this paper, we attempt a proof of RII in $D\geq 4$ using a two-pronged geometric and analytical approach, which we elaborate on now.
Organization of the paper.
The remainder of this article is organized as follows. In Sec. II, we analyze the entropy of AdS–Schwarzschild black holes using the geometric volume, which coincides with the thermodynamic volume in the static case. Using gravitational focusing together with the Sherif–Dunsby conformal rigidity theorem, and independently through a second–variation analysis of the slice functional, we show that the round horizon is the unique entropy maximizer among all hypersurfaces enclosing a fixed thermodynamic volume.
In Sec. III, we extend the discussion to rotating geometries. We first treat Kerr–AdS as an off–shell deformation of the round sphere at fixed thermodynamic volume and show that such a deformation necessarily lowers the entropy relative to the Schwarzschild–AdS case. We then provide an independent thermodynamic argument: since the entropy $S(\mathcal{V},J)$ is strictly concave in $J$ along the thermodynamically stable Kerr–AdS branch, the state with $J=0$ uniquely maximizes entropy at fixed thermodynamic volume $\mathcal{V}$. Together, these two approaches establish that Schwarzschild–AdS is the unique entropy maximizer at fixed thermodynamic volume.
Sec. IV summarizes the main results, delineates the domain of validity of the proof, and outlines several extensions, including modified gravity, non–AdS asymptotics, and quantum or holographic corrections. Appendix A reviews the Sherif–Dunsby rigidity theorem and its relevance to our construction, Appendix B derives the variation of the York boundary term, and Appendix C summarizes the extrinsic geometry of round spheres in Euclidean AdS.
II AdS–Schwarzschild black holes maximize entropy
To begin the analysis, we perform a 1+1+2 split [4] of spacetime $\mathcal{M}$ (since we are only interested in the properties of hypersurface, it is natural to perform a 1+3 or 1+1+2 decomposition of spacetime) of dimension $D=4$ and apply a proper (non-constant) conformal transformation to the metric of the (D1)(D-1)-hypersurface $\Sigma$ of $\mathcal{M}$
$h_{ab}\longrightarrow\Omega^{2}(\mathcal{X})h_{ab}.$ (8)
Here, $\mathcal{X}$ are the angular coordinates, so that the spherical symmetry is broken. For example, one can have $\Phi\equiv\Phi(\theta)$ which explicitly breaks the $SO(3)$ symmetry. In the 1+1+2 decomposition of the spacetime $\mathcal{M}$, the 3-space orthogonal to the timelike unit vector $u^{\mu}$ is further split into a preferred spacelike direction $e^{\mu}$ and its orthogonal 2-space (or “2-sheet”). One introduces the projection tensors [23]
$h_{\mu\nu}=g_{\mu\nu}+u_{\mu}u_{\nu},\qquad N_{\mu\nu}=h_{\mu\nu}-e_{\mu}e_{\nu}\,$,
so that $u^{\mu}u_{\mu}=-1$, $e^{\mu}e_{\mu}=+1$, and $u^{\mu}e_{\mu}=0$. Working on a fixed background (here, Anti-de Sitter) so that $R_{\mu\nu}[g]$ is held fixed by Einstein’s equations, we vary only the intrinsic 3-metric (or induced metric) $h_{\mu\nu}$ via a (proper) conformal transformation. Spatial indices are then raised and lowered with $h_{\mu\nu}$. This point can be understood as follows: Although the induced metric and bulk metric are related by
$h_{ab}=g_{ab}+n_{a}n_{b},$
we hold the bulk metric fixed,
$\delta g_{ab}=0\quad\Longrightarrow\quad\delta R_{ab}[g]=0,$ (9)
and vary only the hypersurface embedding (or its conformal factor). Concretely, deforming the embedding as
$X^{a}(\sigma)\;\longrightarrow\;X^{a}(\sigma)+\Phi(\sigma)\,n^{a}\quad\Longrightarrow\quad n^{a}\;\longrightarrow\;n^{a}+\delta n^{a},$ (10)
induces a change in the induced metric via
$\delta h_{ab}=\delta\bigl(n_{a}n_{b}\bigr)=(\delta n_{a})\,n_{b}+n_{a}\,(\delta n_{b}),$ (11)
even though $\delta g_{ab}=0$. Therefore,
  1. 1.
    Bulk metric fixed: $\delta g_{ab}=0\implies\delta R_{ab}=0.$
  2. 2.
    Hypersurface varied: $X^{a}\to X^{a}+\Phi\,n^{a}$ so that $\delta n^{a}\neq 0$ and hence $\delta h_{ij}\neq 0.$
Thus, we achieve a non‐trivial variation of the induced metric $h_{ij}$ while keeping the ambient Ricci tensor $R_{ab}$ fixed by Einstein’s equations.
We now argue in favor of the round 3-sphere in $\mathcal{M}$ maximizing area or entropy purely on geometric grounds.
II.1 Sherif–Dunsby rigidity and maximal entropy
Consider a compact three–manifold $\Sigma$ of spherical topology in the 1+1+21{+}1{+}2 decomposition of spacetime $\mathcal{M}$. We generate a one–parameter family of non–round slices $\Sigma_{s}$, $s\in(-\varepsilon,\varepsilon)$, by a smooth, volume–preserving normal deformation
$X^{a}\;\mapsto\;X^{a}+\Phi(s)\,n^{a},\qquad\int_{\Sigma}\Phi(s)\,d\mu=0,$ (12)
where $n^{a}$ is the unit normal to $\Sigma$. This produces embedded hypersurfaces with induced metrics $h_{ab}(s)$ while keeping the ambient Einstein geometry fixed ($\delta g_{ab}=0$). Let $\mathcal{T}\equiv\Sigma_{s=\varepsilon}$ denote a generic deformed compact slice.
Geometric input from focusing.
In the 1+1+21{+}1{+}2 formalism, let $e^{a}$ denote the unit vector orthogonal to the spacelike 22–sheets on $\mathcal{T}$, and define the associated sheet–expansion scalar by
$\hat{\theta}\equiv\delta_{\mu}e^{\mu}.$ (13)
The quantity $\hat{\theta}$ measures the logarithmic rate of change of the area element of the transverse 22–sheets along the spacelike congruence generated by $e^{a}$.
The evolution of $\hat{\theta}$ is governed by the Raychaudhuri equation for this spacelike congruence. In an Einstein background with negative cosmological constant,
$R_{ab}=\Lambda g_{ab},\qquad\Lambda<0,$ (14)
the Ricci contribution enters with the contracting sign, so that the curvature term drives the expansion toward smaller values along the congruence (gravitational focusing). Once the sheet expansion becomes non–positive at some stage along the flow, the Raychaudhuri equation ensures that the subsequent evolution favors further contraction. We therefore evaluate the rigidity theorem in the regime where
$\hat{\theta}<0,$ (15)
corresponding to a contracting sheet geometry.
Since $\hat{\theta}$ equals the logarithmic derivative of the 22–area element along $e^{a}$, negative sheet expansion corresponds to a contracting conformal rescaling of the intrinsic metric on $\mathcal{T}$. Consequently there exists a conformally related representative
$\widetilde{h}_{ab}=e^{2\varphi}\,h_{ab},\qquad\varphi<0,$ (16)
which preserves the scalar curvature, $R(\widetilde{h})=R(h)$. Within the 1+1+21{+}1{+}2 decomposition, this identification between negative sheet expansion and a contracting conformal factor is precisely the relation used by Sherif and Dunsby [23].
Sherif–Dunsby rigidity.
By Theorem VII.4 of Sherif and Dunsby [23] (see Appendix A for details of the theorem and its applicability in our proof), a compact 33–manifold with nonnegative scalar curvature that admits a proper (non–homothetic) conformal transformation
$\widetilde{h}_{ab}=e^{2\varphi}h_{ab},\qquad R(\widetilde{h})=R(h),\qquad\varphi<0,$
is necessarily isometric to the round $S^{3}$ (up to a constant scale factor). In our context, the physical input $\hat{\theta}<0$ directly provides the sign condition $\varphi<0$, so all hypotheses of the rigidity theorem are satisfied. Therefore, no volume–preserving normal deformation $\Phi(s)$ of a round $S^{3}$ can yield a distinct extremal geometry: the round $S^{3}$ is rigid.
Geometrically, the Sherif–Dunsby rigidity theorem, combined with gravitational focusing, implies that any non-homothetic, scalar-curvature-preserving (imposed to isolate pure/genuine shape deformations from changes in the intrinsic curvature scale allowing the geometric argument to probe whether a non-round shape can maximize entropy at fixed thermodynamic volume without introducing additional curvature effects), and volume-preserving deformation of the round compact $S^{3}$ in an Einstein background with negative cosmological constant is not admissible. The round $S^{3}$ is therefore the unique stable configuration under these constrained deformations.
Thermodynamically, these constraints correspond precisely to fixing the thermodynamic volume $V$ (via volume-preserving deformations) and fixing the pressure $P$ (by holding the background cosmological constant $\Lambda$ fixed). Stability of the round $S^{3}$ thus translates into the statement that it maximizes the entropy at fixed $(P,V)$, with all conserved charges held fixed. In this sense, the round $S^{3}$ represents the maximally entropic configuration in the appropriate microcanonical ensemble.
Application to black–hole horizons.
A stationary black–hole spacetime admits a smooth spacelike Cauchy slice $\Sigma$ intersecting the horizon cross–section $H\simeq S^{2}$. By compactifying $\Sigma$ (filling in the horizon interior and compactifying at infinity) one obtains a closed 33–manifold $\widehat{\Sigma}\cong S^{3}$. Although $\widehat{\Sigma}$ need not be round in general, the rigidity theorem ensures that the only stable, volume–preserving slice compatible with gravitational focusing is the round $S^{3}$. Within this geometry, the equatorial $S^{2}\subset S^{3}$ uniquely maximizes area at fixed enclosed 33–volume. Consequently, the black–hole horizon that maximizes entropy at fixed thermodynamic volume is the round $S^{2}$—the horizon of the AdS–Schwarzschild black hole.
Thus, the round $S^{2}$ horizon is the unique entropy maximizer in the microcanonical ensemble at fixed $(P,V)$, with $P$ set by the background cosmological constant and $V$ by the volume constraint. This stands in sharp contrast to the flat Euclidean case, where no curvature–induced rigidity exists to enforce such maximality. Having established this geometric extremization principle, we now turn to the area–variation analysis, which reproduces the same result directly from the Euclidean action.
II.2 Effective Functional and Its Variation
We begin with the Euclidean Einstein–Hilbert action supplemented by the Gibbons–Hawking–York (GHY) boundary term in $D$ dimensions in the presence of a cosmological constant,
$\begin{split}I[g]=\;&-\frac{1}{16\pi G}\int_{\mathcal{M}}(R-2\Lambda)\,\sqrt{g}\,d^{D}x\\ &\quad-\;\frac{1}{8\pi G}\int_{\partial\mathcal{M}}K\,\sqrt{\gamma}\,d^{D-1}x,\end{split}$ (17)
where $\gamma_{ab}$ is the induced metric and $K$ the trace of the extrinsic curvature of the boundary. On an Einstein background satisfying $R_{ab}-\tfrac{1}{2}R\,g_{ab}+\Lambda\,g_{ab}=0$, the scalar curvature obeys $R=\tfrac{2D}{D-2}\Lambda\equiv\widetilde{D}\,\Lambda$.
Writing the Euclidean manifold as a foliation by spatial hypersurfaces $\Sigma_{\tau}$ of period $\beta$, the proper spacetime volume of the region between the bolt and the outer boundary may be written as
$V_{\mathcal{M}}=\int_{0}^{\beta}d\tau\int_{\Sigma}N\,\sqrt{h}\,d^{D-1}x=\beta\,V,$
where $N$ is the lapse and $h_{ij}$ the induced metric on $\Sigma$. Since $R=\widetilde{D}\Lambda$ on-shell, the bulk contribution becomes
$I_{\rm bulk}=-\frac{(\widetilde{D}-2)\Lambda}{16\pi G}\int_{\mathcal{M}}\sqrt{g}\,d^{D}x=-\frac{(\widetilde{D}-2)\Lambda\,\beta}{16\pi G}\,V.$ (18)
The variation of the Gibbons–Hawking–York boundary (GHY) term on the same surface gives (see Appendix B)
$\begin{split}&\delta I_{\rm bdy}=-\frac{\delta A}{4G},\end{split}$ (19)
where $A$ is the area of (D2)(D-2)-sphere.
To obtain the total on–shell Euclidean action, the Einstein–Hilbert term is integrated over the region $\mathcal{M}$ bounded internally by the smooth bolt (the Euclidean horizon) and externally by the York boundary $S_{B}$. Applying the Gauss–Codazzi relation,
$R={}^{(D-1)}\!R+K^{2}-K_{ij}K^{ij}-2\nabla_{a}(Kn^{a}-a^{a}),$ (20)
where $^{(D-1)}\!R$ is the intrinsic scalar curvature of the spatial hypersurfaces orthogonal to the unit normal $n^{a}$, $K_{ij}=h_{i}^{\;c}h_{j}^{\;d}\nabla_{c}n_{d}$ is the extrinsic curvature with trace $K=h^{ij}K_{ij}$, and $a^{a}=n^{b}\nabla_{b}n^{a}$ is the acceleration of the normal, the Einstein–Hilbert integrand separates into a bulk curvature part and a total divergence. The latter produces boundary contributions when integrated between the inner and outer limits of $\mathcal{M}$:
$\begin{split}&\int_{\mathcal{M}}\nabla_{a}(Kn^{a}-a^{a})\sqrt{g}\,d^{D}x=\Bigg\{\int_{S_{B}}K\sqrt{\gamma}\,d^{D-1}x\\ &\hskip 142.26378pt-\int_{S_{H}}K\sqrt{\gamma}\,d^{D-1}x\Bigg\},\end{split}$ (21)
where $\gamma_{ij}$ is the induced metric on each constant–$r$ surface, $S_{B}$ denotes the outer York boundary, and $S_{H}$ a small surface of constant proper distance from the bolt. The contribution from $S_{B}$ is regrouped with the non-divergence terms in the Gauss–Codazzi decomposition to yield a volume term, which, when combined with the cosmological constant contribution, forms the bulk volume part of the action. Together with the Gibbons–Hawking–York (GHY) term, this reconstructs the slice functional (24), thereby encapsulating the dependence of the action on the boundary geometry.
By contrast, the inner endpoint at the smooth bolt gives a finite, universal residue obtained by evaluating the above expression in the near–horizon geometry $ds^{2}\simeq\rho^{2}\kappa^{2}d\tau^{2}+d\rho^{2}+r_{h}^{2}d\Omega_{D-2}^{2}$, which yields
$I_{\text{inner-end}}=-\frac{1}{8\pi G}\,\beta\,\kappa\,\mathcal{A}=-\frac{\mathcal{A}}{4G},$ (22)
after imposing the regularity condition $\beta=2\pi/\kappa$. Thus, the familiar horizon area term arises not from an explicit boundary contribution but from the inner endpoint of the Einstein–Hilbert volume integral at the smooth bolt [^1] An equivalent formulation treats the horizon term as arising from a regulated Gibbons–Hawking–York integral on an infinitesimal surface $\mathcal{S}_{\epsilon}$ surrounding the bolt. One introduces a small tubular surface at a proper distance $\rho=\epsilon$ and evaluates $\tfrac{1}{8\pi G}\!\int_{\mathcal{S}_{\epsilon}}\!K\sqrt{\gamma}\,d^{D-1}x$. Near the horizon the extrinsic curvature behaves as $K\simeq 1/\rho$, and the induced metric as $\gamma_{ij}\simeq\mathrm{diag}(\rho^{2}\kappa^{2},r_{h}^{2}\Omega_{ij})$. Integrating over $\tau$ and the sphere gives $I_{\rm GHY}^{(H)}=-\tfrac{\kappa\beta}{8\pi G}\mathcal{A}_{D-2}$. Imposing smoothness of the Euclidean section, $\beta=2\pi/\kappa$, yields the finite limit $I_{\rm GHY}^{(H)}=-\mathcal{A}/4G$. Thus the familiar horizon area term may be regarded either as the finite inner endpoint of the Einstein–Hilbert bulk integral (as in this formulation) or equivalently as the smooth limiting value of the GHY integral on a vanishingly small surface encircling the bolt..
Therefore, the total on–shell Euclidean action may be expressed as
$I_{E}=I_{\rm slice}-\frac{\mathcal{A}}{4G}.$ (23)
where we have defined the slice functional as
$\begin{split}I_{\rm slice}&\propto-\,A[S]\;-\lambda\,V[S],\\ &\qquad\lambda=\frac{(\widetilde{D}-2)\Lambda\,\beta}{16\pi G}.\end{split}$ (24)
up to constants. Here, $S$ is the York boundary (D2)(D-2)-sphere. The cosmological constant $\Lambda$ can be understood as an effective Lagrange multiplier implementing the volume constraint.
Taking a variation of the on–shell action at fixed $(\beta,\Lambda)$ gives
$\delta I_{E}=\delta I_{\rm slice}-\frac{1}{4G}\,\delta\mathcal{A}.$ (25)
Stationarity of the total action, $\delta I_{E}=0$, then implies
$\delta S=\frac{\delta\mathcal{A}}{4G}=\,\delta I_{\rm slice},$ (26)
showing that the horizon entropy variation equals the geometric variation of the York boundary functional.
We now evaluate the first and second variations of the slice functional [^2] We analyze stability using the boundary functional $I_{\rm slice}$ rather than the horizon area directly because the bolt area is only defined on the on–shell, regular geometry and depends non–locally on the full interior solution. In contrast, $I_{\rm slice}$ is a purely boundary object with local first/second variation formulas (Hadamard, Jacobi operator), so its quadratic form on $S_{B}$ cleanly captures stability. At the saddle one has $I_{E}=I_{\rm slice}-\mathcal{A}/4G$, hence $\delta S=\delta\mathcal{A}/4G=\delta I_{\rm slice}$ at fixed $(\beta,\Lambda)$; the entropy (horizon–area) stability is therefore equivalent to the stability of $I_{\rm slice}$. under normal deformations of the boundary surface $S$,
$X^{a}\mapsto X^{a}+\phi\,n^{a},$
where $n^{a}$ is the unit normal and $\phi$ a smooth deformation function on $S$. The first variations of area and volume are given by the Hadamard formulas [^3] The volume functional is defined geometrically as $V[S]=\!\int_{\text{inside }S}\!\!\sqrt{h}\,d^{3}x$. In the static case, this term arises in the Euclidean action as the quantity conjugate to the cosmological constant $\Lambda$, and therefore corresponds to the thermodynamic volume in extended black–hole thermodynamics. Its variation is expressed geometrically by the Hadamard formula, $\delta V=\!\int_{S}\!\phi\,dA$, where $\phi$ denotes the normal deformation of the surface $S$.
$\begin{split}\delta A&=\int_{S}H\,\phi\,dA,\\ \delta V&=\int_{S}\phi\,dA,\end{split}$ (27)
where $H$ is the mean curvature of $S$. We restrict attention to volume–preserving deformations,
$\delta V=0\quad\Longleftrightarrow\quad\int_{S}\phi\,dA=0,$ (28)
so that admissible $\phi$ lie in the zero–mean subspace (removing the $\ell=0$ mode). Equivalently, one may impose this constraint with the effective Lagrange multiplier $\lambda$ defined before by extremizing
$I_{\rm slice}[S]=-A[S]-\lambda\,V[S],$ (29)
whose first variation yields
$\delta I_{\rm slice}=-\int_{S}(H+\lambda)\,\phi\,dA=0\quad\Longrightarrow\quad H=-\lambda=\text{const},$ (30)
i.e., the stationary surface is of constant mean curvature (CMC).
The second variation of area and volume are
$\begin{split}\delta^{2}A&=\int_{S}\!\bigl(|\nabla\phi|^{2}-(|K|^{2}+R_{ab}n^{a}n^{b})\phi^{2}\bigr)\,dA,\\ \delta^{2}V&=-\!\int_{S}H\,\phi^{2}\,dA,\end{split}$ (31)
where $K_{ij}$ is the second fundamental form, $|\nabla\phi|^{2}$ the squared tangential gradient on $S$, and $R_{ab}n^{a}n^{b}$ the ambient Ricci curvature along $n^{a}$. The second variation of the functional is $\delta^{2}I_{\rm slice}=-\delta^{2}A-\lambda\delta^{2}V$. Using the CMC condition $H=-\lambda$, this becomes
$\delta^{2}I_{\rm slice}\big|_{\delta V=0}=\int_{S}\!\bigl(-|\nabla\phi|^{2}+(|K|^{2}+R_{ab}n^{a}n^{b}-H^{2})\phi^{2}\bigr)\,dA.$ (32)
For an Einstein background $R_{ab}=\Lambda g_{ab}$, one has $R_{ab}n^{a}n^{b}=\Lambda$. On a round 22–sphere of radius $R$ embedded in a spatial slice of Euclidean AdS4 with $f(R)=1+R^{2}/l^{2}$ (see Appendix C)
$|K|^{2}=\frac{2f(R)}{R^{2}},\quad H=\frac{2\sqrt{f(R)}}{R},\quad\Lambda=-\frac{3}{l^{2}}.$ (33)
Expanding $\phi$ in spherical harmonics with eigenvalues of the Laplacian $\mu_{\ell}=-\ell(\ell+1)/R^{2}$ (and excluding the $\ell=0$ mode by the volume constraint), the stability of each mode is determined by the sign of
$Q_{\ell}=-\frac{\ell(\ell+1)}{R^{2}}+\Big(\frac{2f(R)}{R^{2}}+\Lambda-H^{2}\Big),\qquad\ell\geq 1.$ (34)
Therefore, the modewise quadratic form for volume preserving ($\delta V=0$) deformation is
$\delta^{2}I_{\rm slice}^{(\ell)}=-\frac{\ell(\ell+1)}{R^{2}}-\frac{2f(R)}{R^{2}}-\frac{3}{l^{2}},$ (35)
where the last term comes from the negative cosmological constant.
Hence, for $\ell\geq 2$ modes, which are volume preserving, true shape deformations of the Laplace-Beltrami spectrum [^4] The $\ell=1$ modes correspond to rigid translations or rotations of the sphere, generated by Killing vectors of the background metric; these are pure diffeomorphisms that leave the intrinsic and extrinsic geometry invariant. Hence, only the first non–trivial volume preserving physical deformation, $\ell=2$, contributes to the shape variation of the horizon or boundary geometry. , we get
$\delta^{2}I_{\rm slice}<0.\qquad(\ell\geq 2)$
This shows that the round Euclidean horizon $S^{2}$ in an AdS space is a local maximum of horizon area $\mathcal{A}$ (or entropy $S$) since $\delta^{2}I_{\rm slice}<0\implies\delta^{2}\mathcal{A}<0$.
It is worth mentioning that in the case of horizons, the area is identified with entropy. Therefore, maximizing area (or entropy) leads to stability, unlike the Euclidean isoperimetric inequality, where minimizing area is related to stability. Although we derived the result for Euclidean AdS space in $D=4$, the analysis applies directly to Euclidean AdS space in $D\geq 4$ since all the results (gravitational focusing, Sherif-Dunsby rigidity, and second area variation) continue to hold. Therefore, the conclusion naturally extends to $D\geq 4$. The extension of the Sherif-Dunsby result can be better understood as follows. Because Obata’s theorem [20] holds on any compact $n$-manifold ($n\geq 2$), we may replace the 3D Yamabe rigidity of Sherif–Dunsby by its $n$-dimensional analogue. Concretely, on a $D$-dimensional spacetime we perform a 1+1+(D2)1+1+(D-2) decomposition and identify the infinitesimal conformal factor with expansion of (D2)(D-2) sheet, $\dot{\phi}=\theta$. Then, via gravitational focusing, we have $\dot{\phi}<0$ everywhere, and Obata’s theorem then forces the deformed metric to be isometric (up to scale) to the round $S^{D-1}$. Hence, the only volume-preserving extremum of the entropy functional is the round sphere.
Analytically, for a round $S^{D-2}$ of radius $R$ embedded in Euclidean AdSD with curvature radius $l$, the second variation of the slice functional for a spherical harmonic mode of degree $\ell$ is
$\begin{split}&\delta^{2}I_{\rm slice}^{(\ell)}=-\Bigg\{\frac{\ell(\ell+D-3)}{R^{2}}+\frac{(D-2)}{R^{2}}+\frac{(2D-3)}{l^{2}}\Bigg\}.\end{split}$ (36)
This equation follows from the eigenvalue spectrum of the Laplace–Beltrami operator on the round sphere $S^{D-2}$ and the extrinsic geometry of a round $S^{D-2}$ in $\mathrm{AdS}_{D}$ (see Appendix C for the D=4D{=}4 case; the generalization is straightforward). Therefore, for the volume–preserving deformation ($\ell\geq 2$)
$\begin{split}&\delta^{2}I_{\rm slice}^{(\ell\geq 2)}\;<\;0\qquad(\text{for $D\geq 4$}).\end{split}$ (37)
Therefore, a round $S^{D-2}$ is a strict local maximum of area (entropy) at fixed volume in all $D\geq 4$.
At this point, it is necessary to analyze another spherically symmetric geometry in GR, the charged (Reissner-Nordström) black hole. It is straightforward to evaluate that adding a charge $Q$ modifies the saddle action as
$I_{\rm saddle}=\frac{1}{4G}(-\mathcal{A}+Q\Psi),$ (38)
where $\Psi$ is the electrostatic potential. The term corresponding to the charge $Q$ appears due to the addition of Maxwell’s action to the Einstein-Hilbert action
$I=I_{EH}-\frac{1}{16\pi G}\int\sqrt{-g}F_{\mu\nu}F^{\mu\nu}d^{D}x.$ (39)
However, the entropy is still given by the usual Bekenstein-Hawking entropy in the fixed $Q$-ensemble
$S_{\rm RN}=\frac{\mathcal{A}}{4G}.$
Concretely, one does a Legendre transform on the bulk Maxwell action by adding the boundary term that fixes the charge, and under these boundary conditions, the total Maxwell variation vanishes. Explicitly, the variation of the bulk Maxwell action is given as
$\delta I_{\rm M}$ $=-\frac{1}{4\pi G}\underbrace{\int_{M}d^{4}x\,\sqrt{-g}\,\nabla_{\mu}F^{\mu\nu}\,\delta A_{\nu}}_{\text{EOM=0}}$
$\hskip 28.45274pt+\frac{1}{4\pi G}\int_{\partial M}d^{3}x\,\sqrt{h}\;n_{\mu}F^{\mu\nu}\,\delta A_{\nu}\,$, (40)
where $h_{ij}$ is the induced metric on $\partial M$ and $n^{\mu}$ its outward normal.
To work in the canonical (fixed-$Q$) ensemble, one adds the boundary term
$I_{\rm bdy}=-\frac{1}{4\pi G}\int_{\partial M}d^{3}x\,\sqrt{h}\;n_{\mu}F^{\mu\nu}A_{\nu}$ (41)
which fixes $Q=\frac{1}{4\pi}\!\displaystyle\int_{S^{2}_{\infty}}n_{\mu}F^{\mu t}\,$. Under these boundary conditions,
$\delta\bigl(I_{\rm M}+I_{\rm bdy}\bigr)=-\frac{1}{4\pi G}\int_{\partial M}d^{3}x\,\sqrt{h}\;A_{\nu}\,\delta\bigl(n_{\mu}F^{\mu\nu}\bigr)=0\,$, (42)
since holding the charged $Q$ fixed at the boundary, implies $\delta(n_{\mu}F^{\mu\nu})=0$. This shows that the Maxwell sector does not contribute to the variation of the on-shell action in the fixed-$Q$ ensemble and one is left with $\delta I_{\rm slice}=-\delta A-\lambda\delta V$ as before.
Evidently, this means that a charged, spherically symmetric black hole also maximizes entropy as required by RII.
III Kerr–AdS and the reverse isoperimetric inequality
In this section, we extend our analysis to include rotation and show that it lowers entropy at fixed thermodynamic volume, establishing AdS–Schwarzschild as the unique entropy maximizer.
III.1 Off–shell deformations and Kerr–AdS
Before we begin the analysis for Kerr–AdS case, it is important to note that we compare entropy $S$ of static and rotating case at a fixed thermodynamic volume $\mathcal{V}$ rather than the geometric volume. For the static case, recall that the volume $V[S]$ appears conjugate to the cosmological constant in the Euclidean action and therefore represents the thermodynamic volume, while for the rotating case, they are not the same. More precisely, the slice functional
$I_{\rm slice}[S]=-A[S]-\lambda\,V[S],$ (43)
arises from the Euclidean action of a static black hole. In the static ensemble the Killing vector is
$\xi_{\rm static}=\partial_{\tau},$
which is hypersurface–orthogonal and generates only the Euclidean time circle of period $\beta=1/T$. The boundary data fixed at infinity are therefore exclusively the temperature $T$ (and pressure through $\Lambda$). In this setting the volume conjugate to the pressure is
$\mathcal{V}=\left(\frac{\partial I_{E}}{\partial P}\right)_{T}=\int_{\text{inside}\,S}\sqrt{h}\,d^{3}x=V[S],$ (44)
so the geometric and thermodynamic volumes coincide. Hence Eq. (43) is the correct off–shell functional for studying volume–preserving deformations of the static Schwarzschild–AdS horizon.
For a rotating spacetime the ensemble is different. The relevant Euclidean Killing vector is
$\xi_{\rm rot}=\partial_{\tau}+\Omega\,\partial_{\phi},$
which is not hypersurface–orthogonal and introduces an additional potential $\Omega$ fixed at the boundary. The rotating ensemble is defined by fixing the boundary data $(\beta,\Omega,\Lambda)$, so that the Euclidean time circle is generated by the Killing vector $\xi_{\rm rot}=\partial_{\tau}+\Omega\,\partial_{\phi}$. In this grand–canonical ensemble the generator of Euclidean time translations is $H_{\rm rot}=M-\Omega J$, and the on–shell Euclidean action evaluates to the corresponding Gibbs potential,
$I_{E}^{\rm Kerr}=\beta(M-\Omega J)-S,$ (45)
where $\Omega$ is the angular velocity measured with respect to a non–rotating frame at infinity. One can note that in extended black–hole thermodynamics the mass $M$ is interpreted as the enthalpy, and the thermodynamic volume is defined unambiguously as the quantity conjugate to the pressure in the first law and Smarr relation (equivalently, the Komar/Iyer–Wald volume). In the rotating grand–canonical ensemble $(T,\Omega,P)$ the appropriate thermodynamic potential is the Gibbs free energy $G=M-TS-\Omega J$, which is computed by the Euclidean on–shell action, $I_{E}=\beta G$. Consequently,
$\mathcal{V}=\left(\frac{\partial G}{\partial P}\right)_{T,\Omega}=\frac{1}{\beta}\left(\frac{\partial I_{E}}{\partial P}\right)_{T,\Omega},$ (46)
so the volume extracted from the Euclidean action necessarily coincides with the thermodynamic volume appearing in the Lorentzian first law and Smarr relation. This identification is standard in extended black–hole thermodynamics; see, for example, Dolan [7] and Cheng et al. [3] (see Appendix D for an explicit computation of Dolan’s result in the rotating case). In Cvetič et al. [5], the thermodynamic volume is defined as the quantity conjugate to the pressure in the enthalpy $M$,
$V=\left(\frac{\partial M}{\partial P}\right)_{S,J}.$ (47)
In our Euclidean formulation, the relevant thermodynamic potential in the rotating grand–canonical ensemble is the Gibbs free energy $G=M-TS-\Omega J$, and the volume is defined as
$\mathcal{V}=\left(\frac{\partial G}{\partial P}\right)_{T,\Omega}.$ (48)
Since $G$ is obtained from $M$ by a Legendre transform with respect to $(S,J)$, both definitions yield the same thermodynamic volume.
Taking an off–shell variation at fixed $(\beta,\Omega)$ gives $\delta I_{E}=\beta(\delta M-\Omega\,\delta J)-\delta S$, so consistency of the variational principle requires the boundary functional to contain the term $\beta\Omega J$. Accordingly, integrating out the bulk Einstein–Hilbert–GHY action subject to these grand–canonical boundary conditions yields the rotating slice functional
$I_{\rm slice}^{\rm rot}[S]=-A[S]-\lambda\,V[S]-\beta\,\Omega\,J[S],$ (49)
which is the natural rotating generalization of the static slice functional (43). Here “off–shell” refers to variations of the York slice $S$ at fixed $(\beta,\Omega,\Lambda)$, with the bulk equations of motion otherwise imposed.
It is convenient to recast Eq. (49) in a static–looking form
$I_{\rm slice}^{\rm rot}[S]=-A[S]-\lambda\,\mathcal{V}[S],$ (50)
by defining the “off–shell” thermodynamic volume functional [^5] We emphasize that the York-surface slice functional $I_{\rm slice}[S]$ is an auxiliary off–shell object introduced solely to implement the variational argument. Off shell, there is no unique or canonical notion of “thermodynamic volume” associated with a generic surface $S$. The physical thermodynamic volume entering black–hole chemistry and the reverse isoperimetric inequality is defined on shell via the extended first law and Smarr relation (equivalently via Komar/Iyer–Wald constructions). In the Euclidean formulation, this volume is extracted as $V=(1/\beta)(\partial I_{E}/\partial P)_{T,\Omega}$ in the appropriate ensemble. Our use of $I_{\rm slice}[S]$ therefore does not introduce a new definition of volume, but provides an off–shell variational framework whose on–shell thermodynamics reproduces the standard result.
$\mathcal{V}[S]=V[S]+\frac{\beta\,\Omega\,J[S]}{\lambda}.$ (51)
This shows that, in the rotating case, the geometric volume appearing in the slice functional must be supplemented by rotation–dependent data; the physical thermodynamic volume is defined only on–shell and, for Kerr–AdS, differs from the naive geometric volume by well–known rotation–dependent contributions. As such, if we use the thermodynamic volume $\mathcal{V}$ for comparison, we can represent the rotating slice functional in the form of static slice functional so as to be able to compare the entropy $S$ at the same thermodynamic volume $\mathcal{V}$ exactly as the conjecture demands. Therefore, the comparison of entropy between static and rotating ensembles must be made using the thermodynamic volume and not the geometric volume.
Our second variation calculation establishes that the round sphere is a strict local entropy maximum at fixed geometric volume (which equals the thermodynamic volume in the static case). The key result is that off–shell deformations of the York boundary functional [^6] The quantity $I_{\rm slice}$ represents the contribution of the outer York boundary to the Euclidean action when evaluated on the regular saddle. It may equivalently be viewed as a geometric functional that captures how the action would change under infinitesimal deformations of the boundary surface at fixed bulk geometry. In this sense, $I_{\rm slice}$ can also be regarded as an “off–shell” functional defined by varying the York surface while keeping the bulk metric fixed, even though in the on–shell decomposition $I_{E}=I_{\rm slice}-\mathcal{A}/4G\,$ it simply represents the boundary part of the total on–shell action. correspond directly to geometric deformations of the horizon area. Specifically, for all spherical harmonic modes $\ell\geq 2$, we find
$\delta^{2}I_{\rm slice}(\ell\geq 2)\;<\;0\quad\implies\quad\delta^{2}\mathcal{A}(\ell\geq 2)\;<\;0\,$, (52)
which shows that the horizon area $\mathcal{A}$ is locally maximized at the round sphere. Moreover, the geometric part of the proof establishes that any finite off–shell deformation away from roundness is not allowed under rigidity and gravitational focusing in AdS space.
These results have direct implications for rotating black holes. Since the space of on–shell solutions forms a subset of the off–shell deformation space,
$\mathcal{D}_{\text{on-shell}}\;\subset\;\mathcal{D}_{\text{off-shell}}\,$, (53)
any entropy reduction established for off–shell deformations necessarily holds for the corresponding on–shell solutions. Turning on rotation corresponds precisely to an $\ell=2$ axisymmetric deformation of the round sphere, whose on–shell realization is the Kerr–AdS spacetime. Therefore, Kerr–AdS horizon inherits the entropy reduction property for all angular momentum $J$ from our general geometric and variational analysis
$S_{\rm Kerr}(\mathcal{V})<S_{\rm Sch}(\mathcal{V}),$ (54)
and do not maximize entropy at fixed thermodynamic volume $\mathcal{V}$—only the spherically symmetric Schwarzschild–AdS solution achieves this maximum.
The fact that Kerr–AdS indeed has strictly lower entropy than Schwarzschild–AdS at fixed thermodynamic volume is confirmed independently in the thermodynamic analysis of $S(\mathcal{V},J)$ in the following subsection, where we work directly in the rotating ensemble.
III.2 Concavity of $S$ in $J$ at fixed thermodynamic volume $\mathcal{V}$
We work on the submanifold of thermodynamic state space with fixed thermodynamic volume $\mathcal{V}$, and introduce the generalized Helmholtz potential
$\begin{split}&\Phi(T,\Omega;V)\;\equiv\;U-TS-\Omega J\;=\;M-P\mathcal{V}-TS-\Omega J,\\ &d\Phi\;=\;-S\,dT\;-\;J\,d\Omega\;-\;P\,d\mathcal{V}.\end{split}$ (55)
Here, $T$ is the temperature, $\Omega$ is the angular velocity, $J$ is the angular momentum, $S$ is the entropy, $P$ is the pressure, and $U=M-P\mathcal{V}$ is the internal energy. The potential $\Phi$ is regarded as a function of the independent variables $(T,\Omega;\mathcal{V})$. A subscript on $\Phi$ indicates a partial derivative with respect to the corresponding variable while holding the others fixed. For example,
$\Phi_{T}\equiv\left(\frac{\partial\Phi}{\partial T}\right)_{\Omega,\mathcal{V}},\quad\Phi_{TT}\equiv\frac{\partial^{2}\Phi}{\partial T^{2}}\Big|_{\Omega,\mathcal{V}},\quad\Phi_{T\Omega}\equiv\frac{\partial^{2}\Phi}{\partial T\,\partial\Omega}\Big|_{\mathcal{V}}.$
On the thermodynamically stable branch, the Hessian of $\Phi$ with respect to $(T,\Omega)$ is negative definite, leading to the stability conditions
$\begin{split}&\Phi_{TT}\;=\;-\,\frac{C_{\Omega,\mathcal{V}}}{T}\;<\;0,\qquad\Phi_{\Omega\Omega}\;=\;-\,\chi_{T,\mathcal{V}}\;<\;0,\\ &\Delta\;\equiv\;\det\!\begin{pmatrix}\Phi_{TT}&\Phi_{T\Omega}\\ \Phi_{T\Omega}&\Phi_{\Omega\Omega}\end{pmatrix}\;>\;0,\end{split}$ (56)
where $C_{\Omega,\mathcal{V}}$ is the specific heat at fixed $(\Omega,\mathcal{V})$ and $\chi_{T,\mathcal{V}}$ is the isothermal moment of inertia at fixed $(T,\mathcal{V})$; both are positive on the stable branch.
Using
$S=-\Phi_{T},\qquad J=-\Phi_{\Omega},$
the map $(T,\Omega)\mapsto(S,J)$ is invertible, since its Jacobian is $\Delta$. One can then express the curvature of $S$ as a function of $J$ at fixed $\mathcal{V}$ entirely in terms of $\Phi$:
$\left(\frac{\partial^{2}S}{\partial J^{2}}\right)_{\mathcal{V}}\;=\;-\frac{\Delta}{\Phi_{TT}\,\Phi_{\Omega\Omega}}\;<\;0.$ (57)
This exact identity shows that $S(\mathcal{V},J)$ is strictly concave in $J$ along the connected thermodynamically stable branch. Combining this global concavity with the fact that $J=0$ is a local maximum of entropy implies [^7] The Reverse Isoperimetric Inequality is a $\mathcal{V}$-only statement: $S(\mathcal{V})\leq S_{\mathrm{Schw}}(\mathcal{V})$. In the thermodynamic subsection, we temporarily resolve the implicit $J$-dependence by writing $S(\mathcal{V},J)$ (equivalently $S_{\mathcal{V}}(J)$) for the entropy of the equilibrium state with fixed thermodynamic volume $\mathcal{V}$ and angular momentum $J$. Here, $J$ merely labels deformations within the fixed-$\mathcal{V}$ Kerr–AdS branch; it is not an independent argument of the RII. Accordingly, statements such as strict concavity $d^{2}S_{\mathcal{V}}/dJ^{2}<0$ (implying $S_{\mathcal{V}}(J)\leq S_{\mathcal{V}}(0)$) serve as consistency checks and do not alter the $\mathcal{V}$-only content of the inequality.
[…]
demonstrating that rotation lowers the entropy at fixed thermodynamic volume $\mathcal{V}$ and that the Schwarzschild–AdS solution ($J=0$) is the global maximizer along the Kerr–AdS family. The above inequality holds for arbitrary $J$ along the entire connected thermodynamically stable branch.
Remarkably, the requirement of thermodynamic stability in the above approach aligns with the observation that super-entropic black holes (which violate reverse isoperimetric inequality) are necessarily thermodynamically unstable in the extended black hole thermodynamics [16].
Scope and validity.
[…]
We emphasize that the assumption (i) should not be regarded as mere technicalities; rather, they precisely delineate the validity regime of the Reverse Isoperimetric Inequality. Outside this regime, violations are known to occur—for example, in configurations with non-compact horizons or along thermodynamically unstable branches such as BTZ and ultra-spinning black holes.
IV Conclusion and Discussion
In this work, we have established a geometric–analytic proof of the reverse isoperimetric inequality (RII) for black holes in Einstein gravity in AdS spacetime for $D\geq 4$. On the one hand, the geometric argument shows, using gravitational focusing together with the Sherif–Dunsby rigidity theorem, that the round horizon $S^{D-2}$ enclosing a fixed thermodynamic volume maximizes the entropy within the class of pure shape deformations considered. On the other hand, the analytic approach demonstrates this using the second variation of entropy: nontrivial volume-preserving ($\ell\geq 2$) perturbations of the round sphere yield $\delta^{2}\mathcal{A}<0$ in AdS space, confirming that the AdS–Schwarzschild horizon is a local maximum of area (entropy) under volume constraints, with the geometric argument providing the corresponding global upgrade within this deformation class. We then extend the analysis to Kerr–AdS black holes and show that they do not maximize the entropy at fixed thermodynamic volume.
Together, these complementary methods not only resolve the long-standing conjecture of RII in extended black hole thermodynamics but also highlight the fundamental role of Einstein’s equations and background curvature in governing entropic extremization.
Several avenues for further investigation naturally arise from our proof:
Effects of Modified Gravity on the RII.
In theories beyond Einstein gravity, such as $f(R)$, Gauss–Bonnet or more general Lovelock gravities, and scalar–tensor models, the gravitational field equations acquire extra curvature‐dependent or scalar‐coupling terms which modify both the Raychaudhuri focusing condition and the form of the Euclidean action functional. In particular, the sheet‐expansion scalar $\theta$ no longer obeys the simple $\theta<0$ condition under volume‐preserving deformations. Analytically, the second variation of $I_{\rm slice}$ acquires extra terms $\delta^{2}I_{\rm higher\text{-}curv}$ coming from variation of the $f(R)$ or Gauss–Bonnet invariants, altering the stability criterion of the extremal hypersurface. A systematic study of these corrections would therefore be required in alternative theories of gravity.
Quantum Corrections.
Incorporating higher-curvature corrections or quantum effects (e.g., via one-loop determinants or entanglement entropy corrections) may modify the geometric rigidity or the second variation functional, potentially leading to refined “quantum RII” bounds.
Holographic Perspectives.
Given the AdS/CFT correspondence, it would be interesting to interpret our RII proof in the dual field theory, perhaps relating maximal horizon entropy at fixed volume to extremal entanglement or energy constraints in the boundary CFT.
Beyond Asymptotic AdS.
Extending the analysis to asymptotically flat or more exotic asymptotics (e.g., Lifshitz, hyperscaling violation) might reveal whether the reverse isoperimetric phenomenon is unique to constant-$\Lambda$ backgrounds or has broader applicability. Moreover, the RII has been advocated to hold for dS black holes in [6]. So, it is interesting to see if the proof can be extended to this case.
Violation in the case of superentropic black holes.
It is known that superentropic black holes violate RII. As part of the proof we presented, this can be traced to their non-compact hypersurface, while the proof requires a compact hypersurface. Nevertheless, a general proof explicitly for non-compact hypersurfaces is an interesting future work.
In summary, our geometric–analytic proof not only proves the reverse isoperimetric conjecture within Einstein gravity but also highlights the deep interplay between spacetime curvature, gravitational focusing, and entropy extremization. We expect these insights to inform future studies of black hole thermodynamics, geometric inequalities in curved manifolds, and the fundamental links between geometry and information in gravitational systems.
Appendix A Conformal Rigidity and Sphericity of Compact Hypersurfaces
In this appendix, we summarize the geometric rigidity result due to Sherif and Dunsby [23] that forms the basis for identifying the unique extremal slice in our proof of the Reverse Isoperimetric Inequality. We then demonstrate that all the hypotheses of this theorem are naturally realized for the hypersurfaces considered in the main text.
Statement of the Theorem
Theorem A.1 (Sherif–Dunsby [23], Theorem VII.4).
Let $(M^{4},g)$ be a spacetime admitting a 1+1+21{+}1{+}2 covariant decomposition, and let $T\hookrightarrow M$ be a compact, smoothly embedded spacelike hypersurface whose induced metric $h_{ab}$ has Ricci tensor
$\mathrm{Ric}_{h}=\alpha\,\mathbf{e}\!\otimes\!\mathbf{e}+\beta\,N,\qquad\alpha\neq\beta,\quad\beta>0,$ (59)
where $\mathbf{e}$ is the unit “radial” direction and $N$ is the projector onto the orthogonal 22–sheets. The condition $\alpha\neq\beta$ means that $(T,h)$ is non–Einstein, i.e. $R_{ab}(h)\neq\lambda h_{ab}$. Suppose further that $T$ admits a proper conformal transformation
$h_{ab}\;\mapsto\;\tilde{h}_{ab}=e^{2\varphi}h_{ab},\qquad\varphi<0,$ (60)
such that the scalar curvature is preserved and nonnegative,
$R(\tilde{h})=R(h)\geq 0,$ (61)
[…]
Here primes denote derivatives along the sheet direction $\mathbf{e}$, and the proper conformal transformation ($\varphi<0$) corresponds to negative sheet expansion in the 1+1+21{+}1{+}2 kinematics.
Application to the Reverse Isoperimetric Inequality
The deformed horizon slice $\mathcal{T}$ considered in Section II.1 satisfies each of the above conditions:
Compactness and positive scalar curvature.
Each hypersurface $\mathcal{T}\simeq S^{3}$ is compact and, by the Yamabe theorem, admits a conformal representative of constant scalar curvature. Since $\mathcal{T}$ lies in an AdS (Einstein) background with negative cosmological constant, the Gauss–Codazzi relation ensures $R(h)>0$ on every spatial slice. Moreover, the event–horizon topology theorem [12] implies that stationary black–hole horizons are of positive Yamabe type, confirming the required curvature sign.
Non–Einstein character.
While the ambient spacetime $(\mathcal{M},g)$ satisfies $R_{ab}=\Lambda g_{ab}$, the deformed slice $\mathcal{T}$ need not: its intrinsic Ricci tensor is not proportional to $h_{ab}$, i.e. it is genuinely non–Einstein. This precisely corresponds to the condition $\alpha\neq\beta$ in the theorem.
Scalar–curvature–preserving conformal deformation.
Within the conformal class $h_{ab}]$, one may choose a representative $\tilde{h}_{ab}=e^{2\varphi}h_{ab}$ satisfying
$R(\tilde{h})=R(h)>0,$
which obeys the elliptic constraint
$4\Delta_{h}\varphi+2|\nabla\varphi|^{2}+R(h)(1-e^{2\varphi})=0.$
This condition freezes the intrinsic scalar curvature, ensuring that variations probe only the extrinsic shape of the hypersurface rather than its internal curvature profile. This allows the geometric argument to probe whether a non-round shape can maximize entropy at fixed thermodynamic volume without introducing additional curvature effects.
Negative conformal factor from gravitational focusing.
In an Einstein background with negative cosmological constant, $R_{ab}=\Lambda g_{ab}$ with $\Lambda<0$. The Raychaudhuri equation for the sheet expansion $\theta=\delta_{\mu}e^{\mu}$ implies gravitational focusing of the spacelike sheet congruence generated by $e^{a}$, driving the congruence toward a contracting regime along the flow. We therefore evaluate the rigidity theorem once the congruence enters the regime
$\theta<0,$ (62)
in which the 1+1+21{+}1{+}2 formalism identifies the geometry with a contracting conformal rescaling of the sheet metric. In particular, the intrinsic metric on $\mathcal{T}$ admits a conformal representative
$\tilde{h}_{ab}=e^{2\varphi}h_{ab},\qquad\varphi<0,$ (63)
which preserves the scalar curvature. Thus the “proper” conformal transformation required by the Sherif–Dunsby theorem arises naturally from the gravitational focusing of the normal congruence.
All the hypotheses of the Sherif–Dunsby theorem are therefore satisfied: $\mathcal{T}$ is compact, of positive scalar curvature, non–Einstein, admits a scalar–curvature–preserving conformal transformation, and possesses a negative conformal factor associated with focusing. Consequently, $(\mathcal{T},h)$ must be isometric to the round $S^{3}$.
This conformal rigidity establishes that no volume– and curvature–preserving deformation of a round three–sphere can yield a distinct extremal geometry. Combined with the second–variation analysis of Section II.2, it follows that the round $S^{3}$ slice is the unique stable extremum of the Euclidean action, and the equatorial $S^{2}$ horizon within it maximizes the entropy (area) at fixed thermodynamic volume. This completes the geometric foundation of the Reverse Isoperimetric Inequality for AdS black holes.
Appendix B Boundary-‐term variation for a general York boundary
We consider the Gibbons–Hawking–York term on a timelike (or spacelike) boundary $\Sigma$,
$I_{\rm bdy}=-\frac{1}{8\pi G}\int_{\Sigma}K\,\sqrt{\gamma}\,d^{D-1}x,$ (64)
where $\gamma_{ij}$ is the induced metric on $\Sigma$ and $K=\gamma^{ij}\nabla_{i}n_{j}$ its extrinsic curvature.
Under an infinitesimal normal deformation
$X^{a}$ $\;\longrightarrow\;X^{a}+\phi(x)\,n^{a},\quad n_{a}n^{a}=\pm 1,$ (65)
the induced volume element varies as
$\delta\sqrt{\gamma}$ $=\tfrac{1}{2}\sqrt{\gamma}\,\gamma^{ij}\,\delta\gamma_{ij}\;=\;\sqrt{\gamma}\,K\,\phi,$ (66)
since $\delta\gamma_{ij}=2\,\phi\,K_{ij}$ and $K=\gamma^{ij}K_{ij}$. Hence
$\delta A_{\Sigma}\equiv\delta\!\int_{\Sigma}\sqrt{\gamma}\,d^{D-1}x=\int_{\Sigma}\delta\sqrt{\gamma}\,d^{D-1}x$
$\hskip 22.76228pt=\int_{\Sigma}K\,\phi\,\sqrt{\gamma}\,d^{D-1}x.$ (67)
On the other hand, one may view $\delta I_{\rm bdy}$ as the difference between two boundaries separated by a thin shell of thickness $\phi$. To first order in $\phi$,
$\delta I_{\rm bdy}$ $=-\frac{1}{8\pi G}\Biggl[\int_{\Sigma^{\prime}}K\,\sqrt{\gamma}\,d^{D-1}x-\int_{\Sigma}K\,\sqrt{\gamma}\,d^{D-1}x\Biggr]$
$=-\frac{1}{8\pi G}\int_{\delta\Sigma}K\,\sqrt{\gamma}\,d^{D-1}x$
$\;\approx\;-\frac{1}{8\pi G}\int_{\Sigma}K\,\phi\,\sqrt{\gamma}\,d^{D-1}x\;=\;-\frac{\beta}{8\pi G}\,\delta A.$
$\hskip 128.0374pt=-\frac{\delta A}{4G}$ (68)
This is the Eq.(19) of the main text.
To understand the thin shell argument rigorously, introduce Gaussian normal coordinates $(u,x^{i})$ in a neighborhood of $\Sigma$:
$\begin{split}ds^{2}&=du^{2}+\gamma_{ij}(u,x)\,dx^{i}\,dx^{j},\\ n^{a}&=\partial_{u},\\ \Sigma:\,u&=0,\quad\Sigma^{\prime}:\,u=\phi(x).\end{split}$ (69)
Then the “shell” $\delta\Sigma$ is the set of points $\{(u,x^{i})\mid 0\leq u\leq\phi(x)\}$, and its volume element is
$dV_{\rm shell}=\sqrt{\det\bigl[\gamma_{ij}(u,x)\bigr]}\;du\;d^{D-1}x.$ (70)
The variation of the GHY boundary action is
$\begin{split}\delta I_{\rm bdy}&=-\frac{1}{8\pi G}\int_{\delta\Sigma}K\,\sqrt{\gamma}\,d^{D-1}x\\ &=-\frac{1}{8\pi G}\int_{\Sigma}d^{D-1}x\int_{0}^{\phi(x)}K(u,x)\,\sqrt{\det[\gamma_{ij}(u,x)]}\;du.\end{split}$ (71)
Expanding to first order in the small displacement $\phi$:
$\begin{split}&K(u,x)\,\sqrt{\det[\gamma_{ij}(u,x)]}=K(0,x)\,\sqrt{\det[\gamma_{ij}(0,x)]}+O(u),\\ &\int_{0}^{\phi(x)}du=\phi(x),\\ &\int_{0}^{\phi(x)}K(u,x)\,\sqrt{\det[\gamma_{ij}(u,x)]}\,du\approx K(0,x)\,\sqrt{\gamma(x)}\,\phi(x).\end{split}$ (72)
Hence, to first order,
$\delta I_{\rm bdy}=-\frac{1}{8\pi G}\int_{\Sigma}K(x)\,\phi(x)\,\sqrt{\gamma(x)}\,d^{D-1}x\;+\;O(\phi^{2}).$ (73)
Appendix C Extrinsic and mean curvature of a 2–sphere embedded in Euclidean AdS4
[…]
$ds^{2}_{H^{3}}=d\rho^{2}+l^{2}\sinh^{2}(\rho/l)\,d\Omega_{2}^{2}.$ (74)
A 2–sphere at fixed $\rho=\rho_{0}$ has a real radius
$R(\rho_{0})=l\sinh(\rho_{0}/l)\;\equiv R,$ (75)
with induced metric
$h_{ab}=R^{2}\,\gamma_{ab},$ (76)
where $\gamma_{ab}$ is the unit $S^{2}$ metric.
The outward pointing unit normal is $n=\partial_{\rho}$. The extrinsic curvature of this 2–sphere is
$K_{ab}=\frac{1}{2}\,\mathcal{L}_{n}h_{ab}=\frac{R^{\prime}(\rho_{0})}{R(\rho_{0})}\,h_{ab},$ (77)
with $R^{\prime}(\rho)=\cosh(\rho/l)$. Thus
$K_{ab}=\frac{1}{l}\coth(\rho_{0}/l)\,h_{ab}=\frac{1}{l}\sqrt{1+\frac{l^{2}}{R^{2}}}\,h_{ab}.$ (78)
The mean curvature is therefore
$H=h^{ab}K_{ab}=\frac{2}{l}\coth(\rho_{0}/l)=\frac{2}{l}\sqrt{1+\frac{l^{2}}{R^{2}}}.$ (79)
Hence, a round 2–sphere of radius $r$ embedded in the hyperbolic 3–space $H^{3}$ is totally umbilic with constant mean curvature $H$ given above. This expression provides the geometric input for the variational analysis in the main text.
Appendix D Rotating generalization of Dolan’s identification $I_{E}=\beta G$
In this Appendix we generalize the standard identification of the Euclidean on–shell gravitational action with a thermodynamic potential (as emphasized for the static case by Dolan [7]) to the rotating grand–canonical ensemble. The key output is the rotating analogue of the familiar relation $I_{E}=\beta F$:
$I_{E}=\beta\,G(T,\Omega,P),\qquad G\equiv M-TS-\Omega J,$ (80)
so that, in extended thermodynamics,
$\left(\frac{\partial I_{E}}{\partial P}\right)_{T,\Omega}=\beta\,V,\qquad V=\left(\frac{\partial G}{\partial P}\right)_{T,\Omega}.$ (81)
This shows that the volume extracted from the Euclidean action is the standard thermodynamic volume appearing in the Lorentzian first law/Smarr relation (and hence in the Komar/Iyer–Wald formulation used in black-hole chemistry).
Rotating grand–canonical boundary data
We consider stationary, asymptotically AdS saddles in Euclidean signature, with boundary data fixed at infinity
$(\beta,\Omega,\Lambda)\qquad\text{equivalently}\qquad(T,\Omega,P),\qquad P\equiv-\frac{\Lambda}{8\pi G}.$ (82)
The Euclidean time circle is generated by the Killing vector
$\xi\;=\;\partial_{\tau}+\Omega\,\partial_{\phi},$ (83)
where $\Omega$ is the angular velocity defined with respect to a non–rotating frame at infinity (the standard convention in extended black-hole thermodynamics). To ensure the metric remains real and positive-definite (Riemannian) in the Euclidean section, we employ the standard analytic continuation $t\to-i\tau$ and $J\to iJ$ (implying $a\to ia$ and $\Omega\to i\Omega_{E}$), as detailed in [7].
The corresponding grand–canonical partition function is
[…]
so semiclassically
$Z(\beta,\Omega,P)\sim e^{-I_{E}[g_{\rm cl}]},$ (85)
with $g_{\rm cl}$ the Euclidean saddle obeying the boundary conditions (82).
Euclidean action and the Hamiltonian generator $H_{\xi}=M-\Omega J$
We start from the regulated Euclidean Einstein action (including the Gibbons–Hawking–York term and the usual AdS counterterms)
$\begin{split}&I_{E}[g]=-\frac{1}{16\pi G}\int_{M}(R-2\Lambda)\sqrt{g}\,d^{4}x-\frac{1}{8\pi G}\int_{\partial M}K\sqrt{\gamma}\,d^{3}x\\ &\qquad+I_{\rm ct}[\gamma].\end{split}$ (86)
For a stationary saddle, the Hamiltonian (ADM) decomposition expresses the on–shell action as a pure boundary term plus a horizon contribution. The boundary term is the conserved charge associated with the Euclidean time generator $\xi$, i.e. the Hamiltonian $H_{\xi}$ that generates translations along (83). In the rotating grand–canonical ensemble this is
$H_{\xi}=M-\Omega J,$ (87)
because $\xi$ combines time translations with a rigid rotation at infinity.
Accordingly, the on–shell Euclidean action takes the universal form
$I_{E}=\beta\,H_{\xi}-S\;=\;\beta\,(M-\Omega J)-S.$ (88)
The entropy term $S$ arises from the horizon (bolt) contribution, exactly as in the static case: it is fixed by regularity of the Euclidean section and equals $A_{H}/(4G)$. Importantly, rotation changes the boundary generator from $M$ to $M-\Omega J$ but does not change the structure of the horizon term.
Equation (88) is the rotating generalization of the standard static identification discussed by Dolan [7].
Identification with Gibbs free energy
Define the Gibbs free energy in the rotating grand–canonical ensemble:
$G(T,\Omega,P)\equiv M-TS-\Omega J.$ (89)
Using $\beta=1/T$, the result (88) immediately gives
$I_{E}=\beta\,G,$ (90)
which establishes (80).
Thermodynamic consequence: $(\partial I_{E}/\partial P)_{T,\Omega}=\beta V$
From the extended first law
$dM=T\,dS+\Omega\,dJ+V\,dP,$ (91)
we obtain
$dG=d(M-TS-\Omega J)=-S\,dT-J\,d\Omega+V\,dP.$ (92)
Multiplying by $\beta$ and using $I_{E}=\beta G$ yields
$dI_{E}=\beta\,dG=\beta\,(V\,dP-S\,dT-J\,d\Omega),$ (93)
and therefore at fixed $(T,\Omega)$,
$\left(\frac{\partial I_{E}}{\partial P}\right)_{T,\Omega}=\beta\,V.$ (94)
This is the precise sense in which the “volume extracted from the Euclidean action” coincides with the standard thermodynamic volume appearing in the Lorentzian first law and Smarr relation. In particular, for Kerr–AdS black holes this $V$ is the same thermodynamic (Komar/Iyer–Wald) volume employed in black-hole chemistry and in the formulation of the reverse isoperimetric inequality.
Explicit Euclidean derivation of the Kerr–AdS thermodynamic volume
In this subsection we explicitly demonstrate, for rotating black holes, that the pressure derivative of the Euclidean on-shell action reproduces the standard Kerr–AdS thermodynamic volume, without assuming a priori the identification $I_{E}=\beta G$. This provides a direct Euclidean confirmation of the Lorentzian extended thermodynamics.
We work in four dimensions and follow the Euclidean action computation of Cheng et al. [3]. The regulated Euclidean action for Kerr–AdS, evaluated on a smooth saddle (with no conical defect), is
$\begin{split}I_{E}(r_{+},a,L;\beta)&=\frac{\beta\,r_{+}}{2G\,\Xi}\left(1+\frac{r_{+}^{2}}{L^{2}}\right)-\frac{\pi(r_{+}^{2}+a^{2})}{G\,\Xi},\\ \Xi&\equiv 1-\frac{a^{2}}{L^{2}},\end{split}$ (95)
where $r_{+}$ is the horizon radius, $a$ the rotation parameter, and $L$ the AdS length. The inverse temperature $\beta$ is fixed by regularity of the Euclidean section.
The rotating grand-canonical ensemble is defined by fixing the boundary data $(\beta,\Omega)$, where the angular velocity measured with respect to a non-rotating frame at infinity is
$\Omega=\frac{a}{r_{+}^{2}+a^{2}}\left(1+\frac{r_{+}^{2}}{L^{2}}\right).$ (96)
The Hawking temperature is
$\begin{split}T_{H}=\frac{r_{+}}{4\pi(r_{+}^{2}+a^{2})}\left(1+\frac{a^{2}}{L^{2}}+\frac{3r_{+}^{2}}{L^{2}}-\frac{a^{2}}{r_{+}^{2}}\right),\end{split}$ (97)
so that $\beta=T_{H}^{-1}$.
We now vary the cosmological constant, or equivalently the AdS length $L$, while holding the ensemble data fixed:
$\delta\beta=0,\qquad\delta\Omega=0.$ (98)
These conditions induce variations of $r_{+}$ and $a$ when $L$ is varied. At linear order the constraints read
$\begin{split}\partial_{r_{+}}T_{H}\,dr_{+}+\partial_{a}T_{H}\,da+\partial_{L}T_{H}\,dL&=0,\\ \partial_{r_{+}}\Omega\,dr_{+}+\partial_{a}\Omega\,da+\partial_{L}\Omega\,dL&=0.\end{split}$ (99)
For later use, the required partial derivatives are
$\begin{split}&\partial_{r_{+}}\Omega=-\frac{2ar_{+}\Xi}{(r_{+}^{2}+a^{2})^{2}},\\ &\partial_{a}\Omega=\frac{(1+r_{+}^{2}/L^{2})(r_{+}^{2}-a^{2})}{(r_{+}^{2}+a^{2})^{2}},\qquad\partial_{L}\Omega=-\frac{2ar_{+}^{2}}{L^{3}(r_{+}^{2}+a^{2})},\end{split}$ (100)
and
$\partial_{a}T_{H}=-\frac{ar_{+}(L^{2}+r_{+}^{2})}{\pi L^{2}(r_{+}^{2}+a^{2})^{2}},\qquad\partial_{L}T_{H}=-\frac{r_{+}(a^{2}+3r_{+}^{2})}{2\pi L^{3}(r_{+}^{2}+a^{2})},$ (101)
with $\partial_{r_{+}}T_{H}$ determined straightforwardly.
Solving (99) yields $dr_{+}/dL$ and $da/dL$, which are then inserted into the total variation of the Euclidean action at fixed $\beta$,
$dI_{E}=\partial_{r_{+}}I_{E}\,dr_{+}+\partial_{a}I_{E}\,da+\partial_{L}I_{E}\,dL.$ (102)
The partial derivatives of $I_{E}$ at fixed $\beta$ are
$\begin{split}&\partial_{r_{+}}I_{E}=\frac{\beta(L^{2}+3r_{+}^{2})-4\pi L^{2}r_{+}}{2GL^{2}\Xi},\\ &\partial_{a}I_{E}=\frac{a(L^{2}+r_{+}^{2})(\beta r_{+}-2\pi L^{2})}{GL^{4}\Xi^{2}},\end{split}$ (103)
$\partial_{L}I_{E}=-\frac{(r_{+}^{2}+a^{2})(\beta r_{+}-2\pi a^{2})}{GL^{3}\Xi^{2}}.$ (104)
We now express the variation in terms of the thermodynamic pressure, defined in four dimensions by
$P=\frac{3}{8\pi GL^{2}},\qquad\frac{dL}{dP}=-\frac{4\pi GL^{3}}{3}.$ (105)
Combining these results and finally imposing the smoothness condition $\beta=T_{H}^{-1}(r_{+},a,L)$, one finds
$\left(\frac{\partial I_{E}}{\partial P}\right)_{\beta,\Omega}=\beta\,V,$ (106)
where
$\begin{split}V=\frac{4\pi r_{+}(r_{+}^{2}+a^{2})}{3\,\Xi}\left[1+\frac{a^{2}}{2r_{+}^{2}}\frac{1+r_{+}^{2}/L^{2}}{\Xi}\right].\end{split}$ (107)
The expression (107) coincides exactly with the Kerr–AdS thermodynamic volume obtained from the Lorentzian extended first law and Smarr relation [5]. Thus, in the rotating grand-canonical ensemble, the Euclidean on-shell action reproduces the correct thermodynamic volume directly through its pressure derivative, without any independent assumption regarding its interpretation as a Gibbs free energy.